7. Anle Corporation has a current stock price of $20.43
and is expected to pay a dividend of $1.20
in one year. Its expected stock price right after paying that dividend is
$22.66.
a. What is​ Anle's equity cost of​ capital?
b. How much of​ Anle's equity cost of capital is expected to be satisfied by dividend yield and how much by capital​ gain?
Answer #1

a.equity cost of capital=(End value-Beginning value+Dividend)/Beginning value  

=(22.66-20.43+1.2)/20.43

=16.79%(Approx)

b.Dividend yield=Dividend/Beginning value  

=(1.2/20.43)

=5.87%(Approx)

Capital gains yield=(End value-Beginning value)/Beginning value  

=(22.66-20.43)/20.43

=10.92%(Approx)